The Magellanic Clouds Newsletter

⋆ ⋆ ⋆

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                               mcnews@astro.uiuc.edu
                http://www.astro.uiuc.edu/projects/mcnews/MCNews.php

No. 78 September 8th, 2004

Letter From The Editor

Dear Colleagues,

The Magellanic Clouds Working Group is pleased to bring you issue 78 of the Magellanic Clouds Newsletter. In this issue, we present 6 submitted abstracts, a update from the University of Basel, and a collection of recent astro-ph listings of potential interest to Magellanic Clouds researchers. As always, this month’s issue is available from the MC News Website in a variety of formats (PDF, HTML, PostScript, and L^{A}TEX).

Best Regards, Bryan Dunne Editor, MC News

Contents

An Update on the University of Basel . . . . . . 1 5 Abstracts of Refereed Papers . . . . . . . . . . . . . 2 1 Conference Proceeding Abstract . . . . . . . . . . 6 Recent astro-ph Listings . . . . . . . . . . . . . . . . . . . . 6

Upcoming Meeting Quicklist

Science with LSST and Other Large Surveys . . Sept 20–22

A Special Update on the University of Basel

Dear Colleagues:

Many of you have written letters in support of the Astronomical Institute of the University of Basel, which together with a number of other subject areas in sciences and humanities was threatened by closure. We know of approximately 200 letters protesting the closure plan, often signed by more than one astronomer. Each of these letters is unique and expresses the specific concerns of its author(s), which greatly increased their impact. One of the concerns that figured prominently in these letters was the breach of tenure. Furthermore, reports about the events in Basel appeared in Nature (2004, Vol.

427, p. 768), Science (2004, Vol. 303, p. 1135), and Physics World (2004, Vol. 17, p. 6). The general public in Switzerland and abroad supported us and our outreach programs with 12,000 signatures on a petition opposing the planned closure.

One of the initial results of these campaigns was that the faculties of natural sciences and of the humanities were given the possibility to work out themselves detailed plans of how the required savings in the affected areas might be accomplished. While these plans mean substantial reductions for us, they also imply that none of our senior staff members will be laid off and that astronomical teaching and research can continue in Basel.

These plans were recently approved by the Council of our university. In addition, the Council took a very important step: It has now explicitly declared that the University of Basel will respect tenure. Although it remains to be seen how the impending reductions will affect us, the overall situation now looks a lot more positive. In particular, I am pleased that astronomy now seems to have a future in Basel. Your participation in the international letter campaign has played a key role in this. Hence, also in the name of my colleagues, our heartfelt thanks for your valuable support!

Sincerely, Eva K. Grebel Professor for Astronomy University of Basel

Abstracts of Refereed Papers

Millimeter Dust Emission from an SMC cold molecular cloud

M. Rubio (1) F.Boulanger (2), F.Rantakyro (3,1), & A. Contursi (4)

(1) Departamento de Astronomia, Universidad de Chile, Santiago, Chile

(2) Institut d’Astrophysique Spatiale, Université Paris Sud, Orsay Cedex, France

(3) European Southern Observatory, Santiago, Chile

(4) Max-Planck-Institute für Extraterrestische Physik, Garching, Germany

We presented SIMBA 1.2 mm continuum observations of SMCB1-1 an SMC molecular cloud known to be distant from massive ionizing stars. The gas mass derived from the SIMBA flux is a factor of ∼ 10 higher than the virial mass previously reported from CO observations. We propose an interpretation where the virial mass underestimates the total mass because the CO emission comes from dense clumps which do not trace the full cloud area and velocity distribution. This work sets an important question for the understanding of star formation in low metallicity gas. Do CO observations grossly underestimate the amount of dense self gravitating clouds where stars can form?

Comments: Accepted Astronomy and Astrophysics Letter e-mail: mrubio@das.uchile.cl

A Statistical Investigation of HI in the Magellanic Bridge

E. Muller (1), S. Stanimirovic (2), E. Rosolowsky (2), L.Staveley-Smith (3)

(1) Arecibo Observatory

(2) University of California, Berkeley

(3) Australia Telescope National Facility

We present results from two statistical analyses applied to an neutral hydrogen (HI) dataset of the nearby tidal bridge in the Magellanic System. Primarily, analyses of the Spatial Power Spectrum suggest that the Magellanic Bridge, historically considered to be a single contiguous feature, may in fact be a projection of two kinematically and morphologically distinct structures. The southern and more obviously turbulent parts appear to show structure organized similarly to the adjacent Small Magellanic Cloud (SMC), while the northern regions are shown to be relatively deficient in large scale power. The extent of modification to the spatial power index by the velocity fluctuations is also highly variant across these parts of the Bridge. We find again that the northern part appears distinct from the southern part and from the SMC, in that the power spectrum is significantly more affected by slower velocity perturbations. We also probe the rate of spectral variation of the HI by measuring the Spectral Correlation Function over selected regions. The results from this analysis highlight a tendency for the HI spectra within the bright parts of the Bridge to have a more persistent correlation in the E-W direction than in the N-S direction. These results are considered to be quantitative evidence for the tidal processes which are thought to have been active throughout the evolution of the Magellanic Bridge.

Comments: Accepted by ApJ WWW: http://xxx.soton.ac.uk/pdf/astro-ph/0408259 e-mail: emuller@naic.edu

New Cepheid Period-Luminosity Relations for the Large Magellanic Cloud: 92 Near-Infrared Light Curves Curves

S. E. Persson, Barry F. Madore, W. Krzeminski, Wendy L. Freedman, M. Roth and D. C. Murphy

Observatories of the Carnegie Institution of Washington, Pasadena, CA, USA

Near-infrared J, H, and Ks photometric measurements of 92 Cepheids in the Large Magellanic Cloud are presented. The stars are spread over the face of the Cloud, their periods range from 3 to 100 days, and their light curves are sampled at an average of 22 phase points per star. The intensityweighted mean magnitudes and colors define period-luminosity-color (PL or PLC) relations whose uncertainties due to differential metal abundance and reddening/extinction effects are minimal. The dispersions in the infrared PL, PLC, and extinction-free period-Wesenheit relations are extremely small, amounting to ¡ 0.10 mag (or 5% in distance).

The orientation of the disk plane of the sample (inclination angle and line of nodes) agrees well with the results of van der Marel & Cioni (2001).

The PL and PLC fits are the best-determined such relationships yet found for any sample of Cepheids, and establish a calibration that can be used to precisely anchor ground- and space-based near-infrared Cepheid data to external galaxies, and also back to Cepheid calibrators in the Galaxy. As an example, we use the Galactic Cepheid calibratio n of Gieren, Fouque & Gomez (1998) to obtain the distance modulus to the centroid of our LMC sample. The true modulus of the LMC is thus determined to be 18.50 +/- 0.05 mag. Currently the dominant source of uncertainty in this number is the scatter in the Galactic calibrator sample.

The PLC fits and dispersions, and the dependence of the PLC upon metal abundance are compared to theoretical versions computed from the work of Alibert et al. (1999). Overall, the agreement is excellent, indicating that at near-infrared wavelengths the slope and dispersion of the PLC depend very weakly upon metal abundance. The shift in the JHK PLC relations is about 0.02 mag for a change in metal abundance from Solar to one-half Solar.

e-mail: barry@ipac.caltech.edu

Constraints on the star formation history of the Large Magellanic Cloud

S. C. Javiel, B. X. Santiago & L. O. Kerber

Universidade Federal do Rio Grande do Sul, Porto Alegre, RS, Brazil

We present the analysis of deep colour-magnitude diagrams (CMDs) of 6 stellar fields in the LMC. The data were obtained using HST/WFPC2 in the F814W (∼I) and F555W (∼V) filters, reaching V_{555} ∼26.5. We discuss and apply a method of correcting CMDs for photometric incompleteness. A method to generate artificial CMDs based on a model star formation history is also developed. This method incorporates photometric error effects, unresolved binaries, reddening and allows use of different forms of the initial mass function and of the SFH itself. We use the Partial Models Method, as presented by Gallart, Aparicio, Carrera and collaborators, for CMD modelling, and include control experiments to prove its validity in search of constraints on the Large Magellanic Cloud star formation history in different regions. Reliable star formation histories for each field are recovered by this method. In all fields, a gap in star formation with t 700 Myr is observed. Field-to-field variations have also been observed. The two fields near the LMC bar present some significant star forming events, having formed both young (t < 1 Gyr) and old (t > 10 Gyr) stars, with a clear gap from 3-6 Gyrs. Two other fields display quite similar SFHs, with increased star formation having taken place at t ∼ 2-3 Gyrs and 6 < t < 10 Gyrs. The remaining two fields present star formation histories closer to uniform, with no clear event of enhanced star formation.

e-mail: basagran@if.ufrgs.br

Comparing the properties of local globular cluster systems: implications for the formation of the Galactic halo

A. D. Mackey & G. F. Gilmore

Institute of Astronomy, Cambridge, UK

We investigate the hypothesis that some fraction of the globular clusters presently observed in the Galactic halo formed in external dwarf galaxies. This is done by means of a detailed comparison between the “old halo”, “young halo,” and “bulge/disk” sub-systems defined by Zinn (1993) and the globular clusters in the LMC, SMC, and Fornax and Sagittarius dwarf spheroidal galaxies. We first use high quality photometry from Hubble Space Telescope images to derive a complete set of uniform measurements of horizontal branch (HB) morphology in the external clusters. We also compile structural and metallicity measurements for these objects and update the database of such measurements for the Galactic globular clusters, including new calculations of HB morphology for 11 objects. Using these data together with recent measurements of globular cluster kinematics and ages we examine the characteristics of the three Galactic cluster sub-systems. Each is quite distinct in terms of their spatial and age distributions, age-metallicity relationships, and typical orbital parameters, although we observe some old halo clusters with ages and orbits more similar to those of young halo objects. In addition, almost all of the Galactic globular clusters with large core radii fall into the young halo sub-system, while the old halo and bulge/disk ensembles are characterized by compact clusters. We demonstrate that the majority of the external globular clusters are essentially indistinguishable from the Galactic young halo objects in terms of HB morphology, but ∼20-30% of external clusters have HB morphologies most similar to the Galactic old halo clusters. We further show that the external clusters have a distribution of core radii which very closely matches that for the young halo objects. The old halo distribution of core radii can be very well represented by a composite distribution formed from ∼83-85% of objects with structures typical of bulge/disk clusters, and ∼15-17% of objects with structures typical of external clusters. Taken together our results fully support the accretion hypothesis. We conclude that all 30 young halo clusters and 15-17% of the old halo clusters (10-12 objects) are of external origin. Based on cluster number counts, we estimate that the Galaxy may have experienced ∼7 merger events with cluster-bearing dwarf-spheroidal-type galaxies during its lifetime, building up ∼45-50% of the mass of the Galactic stellar halo. Finally, we identify a number of old halo objects which have properties characteristic of accreted clusters. Several of the clusters associated with the recently proposed dwarf galaxy in Canis Major fall into this category.

Comments: MNRAS accepted; astro-ph/0408404 WWW: http://www.ast.cam.ac.uk/∼dmackey/halo.html e-mail: dmackey@ast.cam.ac.uk

Abstracts of Non-Refereed Papers

Structure of the Periphery of the Large Magellanic Cloud Revealed by 2MASS

David R. Alves

National Research Council, Goddard Space Flight Center

In order to understand the equilibrium structure of the Large Magellanic Cloud (LMC), I have reexamined the content of the 2MASS database on its periphery. Based on the dereddened colors of the LMC’s red giant branch (RGB) stars, I find an outward radial gradient of decreasing metallicity. The isodensity contours of the RGB stars rule out the existence of a spherical halo on the periphery, and reveal instead an inclined disk. Surprisingly, these density contours also hint that the thickness of the disk traced by metal-weak RGB stars is greater than that traced by the metal-rich RGB. If the stellar population mass-to-light ratio anticorrelates with metallicity, this structure could boost the LMC’s self-microlensing optical depth.

Comments: 4 pages, 3 figures. To appear in the conference proceedings of ”Wide Field Imaging from Space” (published in New Astronomy Reviews), eds. T. McKay, A. Fruchter, and E. Linder. Previously, I submitted address for the online poster to MCNEWS; this is the proceedings paper with more detail. WWW: http://xxx.lanl.gov/abs/astro-ph/0408336 e-mail: alves@lasp680.gsfc.nasa.gov

Recent astro-ph Listings

astro-ph/0409167 Title: The SuperMACHO Microlensing Survey Authors: A. C. Becker, A. Rest, C. Stubbs, G. A. Miknaitis, A. Miceli, R. Covarrubias, S. L. Hawley, C. Aguilera, R. C. Smith, N. B. Suntzeff, K. Olsen, J. L. Prieto, R. Hiriart, A. Garg, D. L. Welch, K. H. Cook, S. Nikolaev, A. Clocchiatti, D. Minniti, S. C. Keller, B. P. Schmidt Comments: To appear in Proceedings of IAU Symposium 225: Impact of Gravitational Lensing on Cosmology, 6 pages

astro-ph/0409165 Title: Clues on post-Asymptotic Giant Branch Evolution and Planetary Nebulae Populations from the Magellanic Clouds Authors: Letizia Stanghellini Comments: 12 pages, 6 figures Invited review to be published in the proceedings of the ESO workshop ”Planetary Nebulae beyond the Milky Way”, Springer-Verlag Fig. 2 in jpg format

astro-ph/0409023 Title: Surface Brightness and Stellar Populations at the Outer Edge of the Large Magellanic Cloud: No Stellar Halo Yet Authors: C. Gallart, P.B. Stetson, E. Hardy, F. Pont, R. Zinn Comments: ApJLett, in press 13 pages including 3 color figures astro-ph/0409016

Title: On the mass of the gravitational lenses in LMC Authors: L. Mancini, S. Calchi Novati, Ph. Jetzer, G. Scarpetta Comments: To appear in the proceedings of the Marcel Grossmann Meeeting X

astro-ph/0408533 Title: The Case of the Off-Center, Levitating Bar in the Large Magellanic Cloud Authors: Dennis Zaritsky (Univ. of Arizona) Comments: 5 emulateapj style pages, accepted for publication in ApJL

astro-ph/0408452 Title: Reconstructing Cepheid Light Curve with Fourier Techniques II: Catalogue of the Fourier Parameters for OGLE LMC Fundamental Mode Cepheids Authors: C. C. Ngeow (UMASS), S. Kanbur (UMASS) Comments: 4 pages, 2 figures. MNRAS submitted

astro-ph/0408336 Title: Structure of the Periphery of the Large Magellanic Cloud Revealed by 2MASS Authors: David R. Alves Comments: 4 pages, 3 figures. To appear in the conference proceedings of ”Wide Field Imaging from Space” (published in New Astronomy Reviews), eds. T. McKay, A. Fruchter, and E. Linder

astro-ph/0408259 Title: A Statistical Investigation of HI in the Magellanic Bridge Authors: E. Muller (Arecibo Observatory), S. Stanimirovic, E. Rosolowsky (UC, Berkeley), L.Staveley-Smith (Australia Telescope National Facility) Comments: 15 pages, Latex. using emulateapj.sty. 11 Figures. Accepted for publication in ApJ

astro-ph/0408235 Title: Disruption time scales of star clusters in different galaxies Authors: Henny J.G.L.M. Lamers, Mark Gieles, Simon F. Portegies Zwart Comments: 7 pages, 4 figures. Accepted for publication in A&A

astro-ph/0408186 Title: The Star Clusters of the Small Magellanic Cloud: Age Distribution Authors: Marc Rafelski (U.Arizona/U.C.L.A), Dennis Zaritsky (U.Arizona) Comments: 11 pages (emulateapj), accepted for publication in the Astronomical Journal. Electronic tables available at: this http URL

astro-ph/0408126 Title: Late-type giant variables in NGC6522, LMC and SMC: How do they differ? Authors: M. Schultheis, I.S. Glass, M.-R. Cioni Comments: 14 pages, 16 figures, A&A accepted, better resolution available from this ftp URL

astro-ph/0408073 Title: The outflowing disks of B[e] supergiants and unclassified B[e] stars Authors: Michaela Kraus (Utrecht University), Marcelo Borges Fernandes (Observatorio Nacional-MCT, Rio de Janeiro) Comments: 4 pages, 3 figures, to appear in: ”The Nature and Evolution of Disks around Hot Stars”, eds.: R. Ignace & K. Gayley

astro-ph/0408039 Title: H-alpha Stacked Images Reveal Large Numbers of PNe in the LMC Authors: Warren A. Reid (Macquarie University, Sydney, Australia), Quentin A. Parker (Macquarie University, Sydney, Australia and Anglo-Australian Observatory, Sydney, Australia) Comments: 6 pages, 3 figures, LaTeX, to be published in Proceedings of the ESO workshop on Planetary Nebulae beyond the Milky Way held at ESO, Garching, May 19-21, 2004

Magellanic Clouds Newsletter No. 78  7

astro-ph/0407516 Title: Australia Telescope Compact Array Survey of Candidate Ultra-Compact and Buried HII Regions in the Magellanic Clouds Authors: Remy Indebetouw, Kelsey E. Johnson, Peter Conti Comments: 53 pp, AJ accepted, version with full-res figures at this http URL

astro-ph/0407506 Title: Chemically inhomogeneous photoionization modelling of the planetary nebula SMC N87 Authors: Y. G. Tsamis, D. Pequignot (Observatoire de Paris - Meudon)

astro-ph/0407485 Title: Period Color and Amplitude Color relations for MACHO project LMC RR Lyraes Authors: S. M. Kanbur, I. Fernando Comments: Sumitted for publication to MNRAS Letters

astro-ph/0407397 Title: Abundance gradient in Local Group galaxies using Asymptotic Giant Branch stars Authors: M.-R.L. Cioni Comments: 5 pages, 4 figures, Contributed talk to the ESO Workshop: ”Planetary Nebulae beyond the Milky Way”, held at ESO-Garching (Germany) on May 19-21, 2004. Conference proceeding will appear in the Springer-Verlag seires of ESO Astrophysics Symposia, Eds. J.R. Walsh & L. Stanghellini

astro-ph/0407350 Title: Variability of X-ray pulsars in a hard energy band observed with INTEGRAL Authors: A.Lutovinov (1), S.Tsygankov (1), M.Revnivtsev (1,2), M.Chernyakova (3), I.Bikmaev (4,5), S.Molkov (1), R.Burenin (1), N.Sakhibullin (4,5) (1 - IKI, Moscow, Russia; 2 - MPA, Garching, Germany; 3 - ISDC, Versoix, Switzerland; 4 - KGU, Kazan, Russia; 5 - AST, Kazan, Russia) Comments: 6 pages, 13 figures, Proceeding of the 5th INTEGRAL Workshop

astro-ph/0407346 Title: Globular clusters and the formation of the outer Galactic halo Authors: Sidney van den Bergh (1), Dougal Mackey (2) ((1) Dominion Astrophysical Observatory, (2) IoA Cambridge) Comments: 7 pages, 7 figures. Accepted for publication in MNRAS

hep-ph/0407342 Title: Neutralino annihilation into gamma-rays in the Milky Way and in external galaxies Authors: N. Fornengo (1,2), L. Pieri (1), S. Scopel (1) ((1) Torino University, (2) INFN, Sezione di Torino) Comments: 17 pages, 9 tables and 14 figures, typeset with ReVTeX4. A Figure and a comment on the recent HESS data have been added. The version on the archive has low-resolution figures. The paper with high resolution figures may be found through this http URL

astro-ph/0407301 Title: Chemical Composition of Planetary Nebulae: The Galaxy and the Magellanic Clouds Authors: W. J. Maciel, R. D. D.Costa, T. E. P. Idiart Comments: 7 pages, 3 figures, uses svmult.cls and graphicx pacakge, to be published in the proceedings of the ESO Workshop ”Planetary Nebulae beyond the Milky Way”, ed. J. R. Walsh, L. Stanghellini, Springer-Verlag

astro-ph/0407281 Title: Dust and Diffuse Interstellar Bands in the z_{a} = 0.524 Absorption System toward AO 0235+164 Authors: V. T. Junkkarinen (1), Ross D. Cohen (1), E. A. Beaver (1), E. M. Burbidge (1), R. W. Lyons (1), G. Madejski (2) ((1) UCSD, (2) Stanford Linear Accelerator Center) Comments: 39 pages, 8 figures, accepted for publication in ApJ

Magellanic Clouds Newsletter No. 78  8